OpenStudy (anonymous):

Find the derivative of f(x) = 4 divided by x at x = 2.

OpenStudy (jtvatsim):

Rewriting question... Find the derivative of \[f(x) = \frac{4}{x}\] at x = 2.

OpenStudy (jtvatsim):

what have you tried so far?

OpenStudy (anonymous):

I thought it would be 2 but 2 isnt part of the answers

OpenStudy (anonymous):

i will make a guess that you have to do this by hand, not by the power rule i.e. you have to compute \[\lim_{x\to 2}\frac{\frac{4}{x}-2}{x-2}\]

OpenStudy (anonymous):

so would I then cancel the x-2 and get 4 as the answer?

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

first off i made an assumption that that is what you needed to compute am i right? there are other ways

OpenStudy (anonymous):

i ned to find the derivative at x=2

OpenStudy (anonymous):

what method do you have at your disposal? by which i mean do you have to work from the definition? and if so, which definition \[f'(a)=\lim_{h\to 0}\frac{f(a+h)-f(a)}{h}\] or \[f'(a)=\lim_{x\to a}\frac{f(x)-f(a)}{x-a}\]

OpenStudy (anonymous):

the second one is what is in my leson i believe

OpenStudy (anonymous):

ok then since \(a=2\) and \(f(a)=f(2)=\frac{4}{2}=2\) you have to compute \[\lim_{x\to 2}\frac{\frac{4}{x}-2}{x-2}\]

OpenStudy (anonymous):

subtract in the numerator first

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

then cancel

OpenStudy (anonymous):

are you guessing? you are actually pretty close

OpenStudy (anonymous):

No 2-2=1 and the bottom is 2-2=1

OpenStudy (anonymous):

you need to do some algebra here if you plug in 2 you get \(\frac{0}{0}\) need to "simplify" that complex fraction first

OpenStudy (anonymous):

btw \(2-2\neq 1\)

OpenStudy (anonymous):

oh wow sory 0

OpenStudy (anonymous):

I was working on another problem and switched numbers

OpenStudy (anonymous):

it is time for algebra that is what is needed

OpenStudy (anonymous):

no wonder my other answer was wrong

OpenStudy (anonymous):

but if it 0/0 where do I go then?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

it is algebra you have a compound fraction, you need to simplify it subtract in the numerator then you will have a nice cancellation then you can plug in 2 for \(x\)

OpenStudy (anonymous):

in other words, first compute \[\frac{4}{x}-2\]

OpenStudy (anonymous):

How am I suposed to subtract them, i dont know how

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

4-2=2/2=1? Is that what you mean @satellite73

OpenStudy (anonymous):

what happened to your variable?

OpenStudy (anonymous):

your answer should be an expression with an \(x\) in it

OpenStudy (anonymous):

I do not know how to do it.

OpenStudy (anonymous):

I switched the x for 2. Isnt that what you said to do? @satellite73

OpenStudy (anonymous):

subtract \[\frac{4}{x}-2\] your answer needs to have an \(x\) in it

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

i don't mean to torture you, if you are not up the the subtraction i will show you have to do it

OpenStudy (anonymous):

I am just really confused, in some questions I would just have to substitute the x and here it is totally different.@satellite73

OpenStudy (anonymous):

no it is algebra is all that is the biggest cause of confusion the actual ideas are not that much \[\frac{4}{5}-2=\frac{4-5\times 2}{5}\]

OpenStudy (anonymous):

similarly \[\frac{4}{x}-2=\frac{4-2x}{x}\]

OpenStudy (anonymous):

Then substitute with 2? So 0/2=0

OpenStudy (anonymous):

Or would you cancel the two x and get 2

OpenStudy (anonymous):

no that was just the numerator

OpenStudy (anonymous):

now you have \[\frac{\frac{4-2x}{x}}{x-2}\]

OpenStudy (anonymous):

which is \[\frac{4-2x}{x(x-2)}\]

OpenStudy (anonymous):

now you can cancel

OpenStudy (anonymous):

I keep geting 0 for the answer. Im literaly losing my mind .Not even kiding :/

OpenStudy (anonymous):

ok lets back up a second and see what the goal is

OpenStudy (anonymous):

you want to replace x by 2, but you cannot

OpenStudy (anonymous):

if you do, you keep getting \(\frac{0}{0}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that means using algebra, you are going to be able to factor and cancel it always works this way

OpenStudy (anonymous):

eventually you will not get \(\frac{0}{0}\) but you will have to cancel first

OpenStudy (anonymous):

since you want to put \(x=2\) and you cannot, what you are gong to cancel is a factor of \(x-2\) in the top and bottom

OpenStudy (anonymous):

that will allow you to replace \(x\) by 2 but we haven't gotten there yet! we need the algebra first

OpenStudy (anonymous):

i will write it out

OpenStudy (anonymous):

Thankyou

OpenStudy (anonymous):

\[\frac{\frac{4}{x}-2}{x-2}=\frac{\frac{4-2x}{x}}{x-2}\] by subtracting in the numerator \[=\frac{4-2x}{x(x-2)}\] by getting rid of the compound fraction \[=\frac{2(2-x)}{x(x-2)}\] by factoring \[=-\frac{2}{x}\] since \(\frac{2-x}{x-2}=1\)

OpenStudy (anonymous):

now you can replace \(x\) by \(2\) in \(-\frac{2}{x}\) and get \(-\frac{2}{2}=-1\)

OpenStudy (anonymous):

two things to note once is in a week you can do this on your head you will say the derivative of \(\frac{4}{x}\) is \(-\frac{4}{x^2}\) and if you plug in 2 you get -1, but of course now you can plug in anything you like, not just 2

OpenStudy (anonymous):

the second thing to not is all the difficultly is with the algebra this is where the rubber hits the road if your algebra is weak you will have a very very hard time the calculus part is very easy, the computations are what cause problems

OpenStudy (anonymous):

Oh okay so whenever I get a problem like that square the bottom and make it negative and I wil get the answer?

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

Then forget what I said

OpenStudy (anonymous):

you have to do what you have to do you will learn short cut rules for finding the derivative, but math does not work like "if i see this i do that" never does

OpenStudy (anonymous):

To be honest I have ben doing math straight for eight hours and Im done

OpenStudy (anonymous):

good! take a break, have a beer forget it for the night

OpenStudy (anonymous):

Haha If only Thanks for your help though!

OpenStudy (anonymous):

yw